OpenStudy (anonymous):

Coral created a painting that was 4 1/3 ft long and 1 1/5 ft wide. What was the area of Coral's painting?

OpenStudy (misty1212):

hi!!

OpenStudy (anonymous):

hi

OpenStudy (misty1212):

area is the product, compute \[\huge 4\tfrac{1}{3}\times 1\tfrac{1}{5}\]

OpenStudy (misty1212):

probably easier if you write them as improper fractions and compute \[\frac{13}{3}\times \frac{6}{5}\]

OpenStudy (misty1212):

cancel first, multiply last

OpenStudy (hhopke):

26 1 -- comes out to 5 -- 5 5

OpenStudy (hhopke):

Or rather, 1 5 -- feet. @misty1212 did a good job explaining it. 5